# BoQ: A Place is Worth a Bag of Learnable Queries

### v1

@@ 4.4 Ablation studies
Number of learnable queries. We present an ablation study in Tab. 5 to provide insights into the impact of varying the number of learnable queries, , on the performance of BoQ. Using ResNet-50 as a backbone and incorporating two BoQ blocks, we observe that performance improves as the number of queries increases. We observe that in less diverse environments, such as Pitts30k, incorporating additional queries yields marginal performance improvements (going from to queries brings as little as R@1 performance gain). However, as demonstrated by the results on AmsterTime, which contains highly diverse images (spanning decades), the model benefits from adding more queries. This is in accordance with the intuition behind BoQ where each query implicitly learns a distinct universal pattern.
@@
Number of BoQ blocks. To assess the influence of the number of BoQ blocks within our architecture, we conduct experiments by varying , the number BoQ blocks utilized. We use a ResNet-18 backbone followed by BoQ blocks, each comprising a bag of learnable queries. We report R@1 performance for each setup. The results, presented in Tab. 6, demonstrate that even when paired with a lightweight backbone like ResNet-18, BoQ remains competitive against state-of-the-art methods such as CosPlace [11] and MixVPR [4], which use ResNet-50 backbone. Best overall performance is observed when employing BoQ blocks.
@@
Backbone architecture. In Tab. 7 we present a comparative performance of BoQ coupled with different ResNet backbone architectures. For each backbone, we crop it at the second last residual block and follow it with two BoQ blocks, each comprising learnable queries. The total number of parameters (trainable + non trainable) is provided. The empirical results showcase that when BoQ is coupled with ResNet-34, a relatively lightweight backbone, it achieves competitive performance compared to NetVLAD, CosPlace and EigenPlaces coupled with ResNet-50 backbone. Interestingly, using ResNet-101, a relatively deeper backbone, we did not achieve better performance than ResNet-50, which could be attributed to memory constraints necessitating a smaller training batch size. Further experiments with different backbones are provided in the Supplementary.
@@
Ablation on the self-attention. Applying self-attention on () brings more stability/performance by adding context between the queries. As shown in Tab. 8, we train a BoQ model with ResNet-18 backbone for epochs. Adding self-attention between the learnable queries brings consistent performance improvement across various benchmarks. Note that the self-attention’s output can be cached during eval/test to avoid recomputation at every test iteration.
| MSLS-val | Pitts30k | Pitts250k | Nordland | |
| BoQ w/out self-att. | ||||
| BoQ with self-att. |
@@ 5 Conclusion
In future work, building upon the concepts discussed in Sec. 3, the spatial information carried in the output of the last encoder, denoted as , presents an opportunity for further enhancement through the application of special reranking strategies.
Acknowledgement: This work has been supported by The Fonds de Recherche du Québec Nature et technologies (FRQNT 254912). We gratefully acknowledge the support of NVIDIA Corporation with the donation of a Quadro RTX 8000 GPU used for our experiments.
@@ 6 More implementation details
Training image size. For our training implementation, we resized the images to pixels to maintain consistency with the training procedures of [4, 3, 2]. This choice of specific resolution directly influences the spatial resolution of the resulting feature maps, which is 20 when using a cropped ResNet backbone [24]. This choice also allows for the use of larger batch sizes.
Inference image size. Considering that many benchmarks contain images of varying sizes and aspect ratios—and often at higher resolutions— we resize the images to a resolution slightly higher than pixels while preserving the original aspect ratio. This approach maintains the integrity of the scenes by keeping the original aspect ratio, and allows the learned queries to interact with bigger feature maps. In Tab. 9 we perform testing at different image sizes (), using a BoQ model trained with images. As we can see, when the images are resized to a height of pixels, there is a consistent improvement in Recall@1 across almost all datasets. This experiment suggests that heights of and may represent an optimal balance between image detail and the model’s capacity to extract and utilize informative features. Note that the performance gains from resizing to these heights are marginal compared to the baseline size of pixels.
@@
7 Vision Transformer backbones
For this experiment, we trained our technique using two distinct Vision Transformer backbones, each coupled with a single BoQ block, and trained on GSV-Cities dataset [3]. The first backbone, ViT [19], pretrained base variant with M parameters, of which we froze all but the last two blocks to allow for fine-tuning. This model is designed to process fixed input image sizes of . The second backbone, DinoV2 [37], also with M parameters, underwent a similar process of freezing, leaving the final two blocks unfrozen for training. DinoV2+BoQ was trained with images resized to and tested with images resized to .
The results presented in Tab. 10 indicate the impact of Vision Transformer backbone on BoQ model’s performance. For ViT+BoQ, performance was possibly hindered by ViT’s fixed input size of leading to performance degradation. This is particularly noticeable on MSLS, SPED and AmsterTime datasets, where the Recall@1 performance is notably lower than ResNet-50 + BoQ. In contrast, DinoV2+BoQ pushes the boundaries to achieve new state-of-the-art results. The increased Recall@1 scores across all benchmarks are substantial, especially on MSLS, Pittsbug, SPED and AmsterTime. These substantial gains underscore DinoV2’s capability in image feature extraction, thereby enhancing BoQ model’s performance.
| MSLS | Pitts250k | Pitts30k | SPED | AmsterTime | Eynsham | SVOX | |
| ViT+BoQ | |||||||
| DinoV2+BoQ |
@@ 8 Interpretability of the learned queries
The cross-attention mechanisms in our BoQ model have been instrumental in achieving fine-grained feature discrimination, as demonstrated by Fig. 5, Fig. 6 and Fig. 7. These figures provide a visualization of the learned queries’ attention patterns across diverse urban scenes and under various environmental conditions.
Fig. 5 demonstrates the model’s temporal robustness, displaying consistent attention across images of the same location captured at different times. The learned queries reliably focus on distinctive features, such as buildings, foliage, and poles, despite variations in viewpoint, lighting, and weather conditions.
Moving objects within a scene often pose a challenge for VPR techniques. Nonetheless, as shown in Fig. 6 our learned queries focus their attention towards static elements of the environment, avoiding moving objects like vehicles.
Fig. 7 underscores the specialization of the learned queries, showcasing their selective focus on relevant features, such as vegetation and buildings. This selective attention is indicative of our model’s ability to interpret complex visual information within the environment.

### v2

@@ 4.4 Ablation studies
Number of learnable queries. We conduct an ablation study to investigate the impact of varying the number of learnable queries, , on the performance of BoQ. We use a ResNet-50 backbone and two BoQ blocks. The results are presented in Tab.Ā 5. We observe that performance improves as the number of queries increases. However, the benefits of additional queries vary across datasets. In less diverse environments, such as Pitts30k, adding more queries yields marginal performance improvements (increasing from to queries results in a mere R@1 performance gain). In contrast, the results on AmsterTime, which features highly diverse images spanning decades, demonstrate that the model benefits significantly from additional queries. This aligns with the underlying intuition of BoQ, where each query implicitly learns a distinct universal pattern.
@@
| MSLS-val | Pitts30k | Pitts250k | Nordland | |
| BoQ w/out self-att. | ||||
| BoQ with self-att. |
@@
Number of BoQ blocks. To assess the influence of the number of BoQ blocks within our architecture, we conduct experiments by varying , the number BoQ blocks utilized. We use a ResNet-18 backbone followed by BoQ blocks, each comprising a bag of learnable queries. We report R@1 performance for each setup. The results, presented in Tab.Ā 6, demonstrate that even when paired with a lightweight backbone like ResNet-18, BoQ remains competitive against state-of-the-art methods such as CosPlaceĀ [11] and MixVPRĀ [4], which use ResNet-50 backbone. Best overall performance is observed when employing BoQ blocks.
Ablation on the self-attention. Applying self-attention on () brings more stability/performance by adding context between the queries. As shown in Tab.Ā 7, we train a BoQ model with ResNet-18 backbone for epochs. Adding self-attention between the learnable queries brings consistent performance improvement across various benchmarks. Note that the self-attentionās output can be cached during eval/test to avoid recomputation at every test iteration.
Backbone architecture. In Tab.Ā 8 we present a comparative performance of BoQ coupled with different ResNet backbone architectures. For each backbone, we crop it at the second last residual block and follow it with two BoQ blocks, each comprising learnable queries. The total number of parameters (trainable + non trainable) is provided. The empirical results showcase that when BoQ is coupled with ResNet-34, a relatively lightweight backbone, it achieves competitive performance compared to NetVLAD, CosPlace and EigenPlaces coupled with ResNet-50 backbone. Interestingly, using ResNet-101, a relatively deeper backbone, we did not achieve better performance than ResNet-50, which could be attributed to memory constraints necessitating a smaller training batch size.
4.5 Vision Transformer Backbones
All previous experiments in this paper were conducted using a ResNetĀ [25] backbone to ensure fairness against existing state-of-the-art methods. Recently, vision foundation models such as DINOv2Ā [38] have been introduced and rapidly adopted as the de facto backbone for various computer vision tasks. In this experiment, we evaluate the performance of BoQ when using a DINOv2 backbone. We employ the pre-trained base variant with 86M parameters, freezing all but the last two blocks to enable fine-tuning. We train on GSV-CitiesĀ [3] with images resized to and perform testing with images resized to . The results presented in Tab.Ā 9, demonstrate that DINOv2 further enhances the performance of BoQ, achieving new all-time high performance on all benchmarks.
| Test Dataset | DinoV2-BoQ | ResNet50-BoQ | ||||
| R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | |
| MSLS-valĀ [53] | ||||||
| Pitts30k-testĀ [46] | ||||||
| Pitts250k-testĀ [46] | ||||||
| Nordland**Ā [44] | ||||||
| Nordland*Ā [56] | ||||||
| SPEDĀ [56] | ||||||
| AmesterTimeĀ [54] | ||||||
| EynshamĀ [18] | ||||||
| St-LuciaĀ [36] | ||||||
| SVOX-nightĀ [10] | ||||||
| SVOX-sunĀ [10] | ||||||
| SVOX-rainĀ [10] | ||||||
| SVOX-snowĀ [10] | ||||||
| SVOX-overcastĀ [10] | ||||||
| SVOX-allĀ [10] | ||||||
| Tokyo 24/7Ā [47] | ||||||
| MSLS-ChallengeĀ [53] | ā | ā | ā | |||
| SF-LandmarkĀ [16] | ā | ā | ā | |||
@@ 5 Conclusion
In future work, building upon the concepts discussed in Sec.Ā 3, the spatial information carried in the output of the last encoder, denoted as , presents an opportunity for further enhancement through the application of special reranking strategies.
Acknowledgement: This work has been supported by The Fonds de Recherche du QuƩbec Nature et technologies (FRQNT 254912). We gratefully acknowledge the support of NVIDIA Corporation with the donation of a Quadro RTX 8000 GPU used for our experiments.
@@ 6 More implementation details
Training image size. For our training implementation, we resized the images to pixels to maintain consistency with the training procedures of [4, 3, 2]. This choice of specific resolution directly influences the spatial resolution of the resulting feature maps, which is 20 when using a cropped ResNet backboneĀ [25]. This choice also allows for the use of larger batch sizes.
Inference image size. Considering that many benchmarks contain images of varying sizes and aspect ratiosāand often at higher resolutionsā we resize the images to a resolution slightly higher than pixels while preserving the original aspect ratio. This approach maintains the integrity of the scenes by keeping the original aspect ratio, and allows the learned queries to interact with bigger feature maps. In Tab.Ā 10 we perform testing at different image sizes (), using a BoQ model trained with images. As we can see, when the images are resized to a height of pixels, there is a consistent improvement in Recall@1 across almost all datasets. This experiment suggests that heights of and may represent an optimal balance between image detail and the modelās capacity to extract and utilize informative features. Note that the performance gains from resizing to these heights are marginal compared to the baseline size of pixels.
@@ 7 Interpretability of the learned queries
The cross-attention mechanisms in our BoQ model have been instrumental in achieving fine-grained feature discrimination, as demonstrated by Fig.Ā 5, Fig.Ā 6 and Fig.Ā 7. These figures provide a visualization of the learned queriesā attention patterns across diverse urban scenes and under various environmental conditions.
Fig.Ā 5 demonstrates the modelās temporal robustness, displaying consistent attention across images of the same location captured at different times. The learned queries reliably focus on distinctive features, such as buildings, foliage, and poles, despite variations in viewpoint, lighting, and weather conditions.
Moving objects within a scene often pose a challenge for VPR techniques. Nonetheless, as shown in Fig.Ā 6 our learned queries focus their attention towards static elements of the environment, avoiding moving objects like vehicles.
Fig.Ā 7 underscores the specialization of the learned queries, showcasing their selective focus on relevant features, such as vegetation and buildings. This selective attention is indicative of our modelās ability to interpret complex visual information within the environment.